# Distinguish the EMRI and B-EMRI system by gravitational waves

### v1

@@ Abstract
Binary stars are as frequency as single stars in the Universe, and at least 70% of the massive stars located in our Galaxy belong to a binary system. For the first time we produce the gravitational waveforms for the extreme mass ratio inspiral systems (EMRIs) of binary stars moving around central supermassive black hole (SBH).We calculate the orbits of such systems via the Hamilton-Jacobi approach. To improve accuracy we adopt the quadrupole-octupole expression of gravitational wave (GW) and consider the contribution of radiation reaction. Compared to the waveforms of single star, double or multiple spikes emerge in the waveforms of binary stars. We calculate the mismatch of the EMRI waveforms of binary stars and those of single stars, and find the mismatch increases as the binary internal semi-latus rectum $\tilde{p}$ decreases. According to the distinguishable criterion of two signals, it’s inferred that the two types of waveforms are distinguishable for small enough $\tilde{p}$. Further, we demonstrate that the most favourable position to distinguish the waveforms of EMRIs of single star from the waveforms of EMRIs of binary stars is that revolutions of the binary stars and the central SBH takes the same direction, i.e., the rotation axis of the binary is parallel to that of the central SBH.
@@ 1 Introduction
Black hole (BH) may be the most thoroughly-studied object before its discovery in the history of science. Recently, the Event Horizon Telescope (EHT) collaboration announced the images of the SBHs M87* and SgrA*[1, 2], these discoveries together with the detections of GWs[3, 4] confirm the existence of black holes, and launch a new era of testing gravity in strong-field regime and far beyond the scale of solar system.
@@
The goal of next generation GW detector—Laser Interferometer Space Antenna (LISA)—is designed to be most sensitive to detect low frequency GW signals in the band about $110^{-4}-$ 1 Hz. Taiji and Tianqin are designed to probe GWs at similar band. EMRIs are the sources generating GWs in this band. EMRIs are composed of central SBH and stellar-mass compact objects moving around the SBH. The multi-body interaction in the cusp of the stellar population surrounding the SBH may scatter the compact objects into the orbits close the central SBH and form the EMRIs[5, 6]. The GWs generated by EMRIs carry the information both of the background BH and the stellar-mass compact objects, so detections of GWs will help to learn more about the the strong-field character of the BH spacetime, the multipole structure of the BH, the structures of the small body compact objects, and the information of the underlying theory governing the way in which our universe works.
@@
Analysis of the data released by GW detectors replies heavily on the matched filtering technique, which makes correlations between data and template waveforms, thus data analysis require establishing the template waveforms of EMRIs.
@@
The waveforms generated by motions of single stellar-mass compact object in BH spacetime have been studied intensively[7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], including probe fundamental fields[25, 26], modified gravity[27, 28, 29, 30, 31, 32], extra dimensions[33], electric charge and internal structure of small body objects[34, 36, 35, 37, 38, 39, 40], etc. Binary-star systems are important astrophysical objects which are constituted of binary stars revolving around each other. The amount of binary-star systems is huge in the Galaxy, it’s estimated that the amount of binary-star systems is no less than that of single star. The vast majority of stars are located in a binary system. It is estimated that more than 70% of all massive stars may exist in systems with two or three stars. In active galactic nuclei, due to the rich forms of interactions between SBH, the gaseous accretion disks and the stellar mass compact objects, the binary may be formed and scattered onto the orbits close the central SBH to form the systems which we call B-EMRIs. The binary stars moving in central SBH constitute a significant typical EMIRs. Thus, it is quite sensible to consider the waveforms produced by motions of binary stars surrounding a SBH other than the case of single star. The GW generated by each of the binary star overlaps, so the waveforms generated by this type of EMRIs are distinguishable from that generated by single star. In this paper, we just consider the case that the amplitudes of the GW generated by internal motion of binary stars are small compared to that generated by external motion of the mass center, so it’s expected small perturbations would appear when compared to the waveforms of single star. It’s really worthwhile to build the template waveforms of such systems for future data analysis to distinguish the waveforms of the binary-star systems and that of single star.
@@
Since the traditional method of calculating gravitational waveforms—solving Teukolsky equations—cost a lot of time and computational resources[41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], while realistic data analysis need to handle a large number of waveforms, thus more practical methods are needed. In this paper, we adopt the “numerical kludge” (NK) method which generates waveforms more conveniently and quickly but captures the principle feature of true waveforms[54]. To improve accuracy, aside from the mass quadrupole we take into account the mass octupole and current quadrupole moments of the source to solve the wave equations. It’s justified the NK method does excellently at approximating the true gravitational waveforms so this method has already been used for scoping out data analysis of LISA. For the radiation reaction, we consider the adiabatic evolution of orbit parameters: the timescale of evolution of the orbit parameters is much larger than that of orbit cycles such that the motion of mass point in BH spacetime is nearly geodesic. Further, since the amplitude of GWs generated by internal motions are small compared to that generated by external motions of the mass center, to simplify the question we only consider the radiation reaction due to the GW emissions generated by the motion of mass center. We take the hybrid scheme that combines Tagoshi’s[46], Shibata’s[53] and Ryan’s fluxes[55, 56], which performs well for generic circular-inclined orbits when compared to Teukolsky-based results[57].
@@ 2 Numerical Kludge Waveforms
The binary are supposed to revolve around each other along elliptic orbit. We introduce the relative coordinates $\vec{\tilde{r}}=\vec{r}_{1}-\vec{r}_{2}$, with $\tilde{r}=\frac{\tilde{p}}{1+\tilde{e}\cos\tilde{\beta}}$, then the coordinates of the binary with masses $m_{1}$ and $m_{2}$ can be expressed with coordinates of mass center and relative coordinates as $\vec{r}_{1}=\vec{r}_{C}+\vec{\tilde{r}}_{1},\vec{r}_{2}=\vec{r}_{C}+\vec{\tilde{r}}_{2}$, where $\vec{\tilde{r}}_{1}=\frac{m_{1}}{m_{1}+m_{2}}\vec{\tilde{r}},\;\vec{\tilde{r}}_{2}=-\frac{m_{2}}{m_{1}+m_{2}}\vec{\tilde{r}}$. We use the coordinates and parameters with tilde to describe the internal motion of binary stars around each other. The spin axis of the central SBH is along $z_{0}$, as displayed in Fig.1, $\theta_{0}$ is the angle between $z_{0}$ and $z$ which is the normal direction of the surface of binary stars’ internal elliptic orbit. $\phi_{0}$ is the angle between $x_{0}$ and the node line which is the intersection line between the surface of binary stars’ internal orbit and the horizonal surface.
Figure 1: The geometry of binary stars orbiting around the central SBH.
@@
We know Kerr BH admits two obvious Killing vectors $\partial_{t}$ and $\partial_{\phi}$, which correspond to two conserved charges, energy $E$ and angular momentum $L_{z}$, thus for the mass point with mass $\mu$, we suppose the action function to be of the form
@@
which implies $p_{r}=dS_{r}/dr$ and $p_{\theta}=dS_{\theta}/d\theta$. Inserting (3) into the Hamilton-Jacobi equation (2) and separating variables we have
@@
Figure 2: $h_{+}$ waveforms of single and binary stars. For the binary the masses are taken as $m_{1}=M_{\odot}$ and $m_{2}=1.1M_{\odot}$. The mass of central SBH is set to be $10^{6}M_{\odot}$. The (external) orbit parameters of the upper two panels are set to be $p(0)=20M,e(0)=0.3,a=0.9$, the internal orbit parameters of the upper right panel are set to be $\tilde{p}=5M,\tilde{e}=0.2$. The (external) orbit parameters of the lower two panels are set to be $p(0)=10M,e(0)=0.3,a=0.9$, the internal orbit parameters of the lower right panel are set to be $\tilde{p}=3M,\tilde{e}=0.3$. One sees that extra burrs appear in the waveforms of B-EMRI compared to the single star revolving around a SBH.
@@
where $\gamma=E\left[\left(r^{2}+a^{2}\right)^{2}/\Delta-a^{2}\right]-\frac{2MraL_{z}}{\Delta}$. Solving these equations allow us to know the positions of mass points at any time, which prepares us for calculating GWs in the following.
Figure 3: The orbits of compact objects generating the waveforms in Fig.2. For the binary we just display the orbit of one object of the two. The alignment of the panels correspond to the panels of waveforms in Fig.2.
@@
Figure 4: Comparisons between the EMRI $h_{+}$ waveforms of single star (the upper left panel) and that of the binary with distinct semi-latus rectum $\tilde{p}$ (the left five panels). The binary masses are fixed to be $m_{1}=M_{\odot}$ and $m_{2}=1.1M_{\odot}$, and mass of central SBH is $10^{6}M_{\odot}$. The (external) orbit parameters are set to be $p(0)=10M,e(0)=0.3,a=0.9$, the internal orbit parameters of the binary stars are set to be $\tilde{e}=0.2,\theta_{0}=\frac{\pi}{3}$ and $\tilde{p}=5M,4M,3M,2M,M$ respectively.
In the NK prescription, a flat-space trajectory which is “equivalent” to a geodesic in Kerr spacetime is reconstructed by projecting the Boyer-Lindquist coordinates onto a fictitious spherical polar grid, the corresponding Cartesian coordinates are defined as
$x_{0}$ $=r\sin\theta\cos\phi,$ (16)
$y_{0}$ $=r\sin\theta\sin\phi,$
$z_{0}$ $=r\cos\theta.$
Then a particle moving along geodesic path in Kerr spacetime can be viewed as the particle moving along curved path in flat-space, thus it’s convenient for us to calculate GWs with Cartesian coordinates.
The expression (14) are valid for a general extended source in flat-space. For a single mass point with mass $\mu$ the energy-momentum tensor in flat spacetime is given by
@@
For binary stars we know the energy-momentum tensor is the sum of two such terms. The momenta (15) now can be simplified to
@@
Substituting (18) into (14) one then obtains the GW generated by mass point.
@@
where the angles $\{\Theta,\Phi\}$ denote the observation point¡¯s latitude and azimuth respectively. After TT projection the waveform is given by
@@
Figure 5: Comparisons of amplitude of GWs between pure external and pure internal motions of binary. The upper left pannel displays the waveforms generated by external motions of center of mass revolving around SBH, the left three pannels display the waveforms generated by pure internal motions of the binary revolving around each other. The external orbit parameters are taken to be identical with that in Fig.4. The semi-latus rectum of the internal orbits are set to be $\tilde{p}=2M,0.5M,0.2M$ respectively.
@@
where $q=a/M$. We defer the expressions of the coefficients $g_{1}(e)\sim g_{14}(e)$ to the Appendix. This expression of fluxes are the results of combining Tagoshi’s[46], Shibata’s [53] and Ryan’s fluxes [55, 56]. Numerical analysis indicates that these results ameliorate the unphysical inspiral properties for nearly circular and polar orbits, such as the inspiral moves rapidly away from circularity for the trajectories near zero eccentricity $e=0$, and there is discontinuity between trajectories with $\iota$ slightly less than $90^{\circ}$ and those slightly above $90^{\circ}$, etc., thus the fluxes (22) produce more physical reasonable results than previous ones. From (22), we obtain the evolution of eccentricity, semi-latus rectum and inclination angle
@@
$\dot{\iota}$ $=\frac{\partial\iota}{\partial E}\dot{E}+\frac{\partial\iota}{\partial Lz}\dot{Lz}+\frac{\partial\iota}{\partial Q}\dot{Q}.$
@@
Figure 6: Waveforms of single compact object with mass equals to $18M_{\odot}$ and binary with masses $8M_{\odot}$ and $10M_{\odot}$ revolving around central SBH with mass $10^{6}M_{\odot}$. The (external) orbit parameters are set to be $p(0)=10M,e(0)=0.3$, the internal motion parameters are set to be $\tilde{p}=2M$ and $\tilde{e}=0.2$.

### v2

# Gravitational waves excited by binary stars orbiting around a supermassive black hole

@@ Abstract
Binary stars are as frequency as single stars in the Universe, and at least 70% of the massive stars located in our Galaxy belong to a binary system. We produce the gravitational waveforms for the extreme mass ratio inspiral systems (EMRIs) of binary stars moving around central supermassive black hole (SBH), or called B-EMRIs.We calculate the orbits of such systems via the Hamilton-Jacobi approach. To improve accuracy we adopt the quadrupole-octupole expression of gravitational wave (GW) and study the contribution of radiation reaction. Compared to the waveforms of EMRIs, there are higher frequency oscillations superposed on the waveforms of B-EMRIs. We perform frequency spectrum analysis of the GW waveforms, and find that higher frequency signals give their prominency in the waveforms of B-EMRIs. We study the binary compact objects with different masses and inclination angles of the binary internal orbit plane, and obtain the frequency spectrum variation with respect to the parameters. To obtain more precise result for future observation of GWs from space-based detector, we take into account gravito-electromagnetic (GEM) force for the binary orbit on the equatorial plane, and compare the waveforms with that of EMRIs without GEM force. The result of mismatch shows that the waveforms of B-EMRIs are credibly distinguishable by the space-based GW detectors when GEM force is considered.
@@ 1 Introduction
Black hole (BH) may be the most thoroughly-studied object before its discovery in the history of science. Recently, the Event Horizon Telescope (EHT) collaboration announced the images of the SBHs M87* and SgrA*[1, 2]. Such images together with the detections of GWs[3, 4] confirm the existence of black holes, and launch a new era of testing gravity in strong-field regime and far beyond the scale of solar system.
@@
The goal of next generation space-based GW detectors—Taiji, Tianqin, and Laser Interferometer Space Antenna (LISA)—is designed to be most sensitive to detect low frequency GW signals in the band about $110^{-4}-$ 1 Hz. EMRIs are the sources generating GWs in this band. EMRIs are composed of central SBH and stellar-mass compact objects moving around the SBH. The multi-body interaction in the cusp of the stellar population surrounding the SBH may scatter the compact objects into the orbits close the central SBH and form the EMRIs[5, 6]. The GWs generated by EMRIs carry the information both of the background BH and the stellar-mass compact objects, so detections of GWs will help to learn more about the the strong-field character of the BH spacetime, the multipole structure of the BH, the structures of the small body compact objects, and the information of the underlying theory governing the way in which our universe works.
@@
Analysis of the data released by GW detectors replies heavily on the matched filtering technique, which makes correlations between data and template waveforms, thus establishing template waveforms of EMRIs is critical for data analysis.
@@
The waveforms generated by motions of single stellar-mass compact object in BH spacetime have been studied intensively[7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], including probe of fundamental fields[25, 26], modified gravity[27, 28, 29, 30, 31, 32], extra dimensions[33], electric charge and internal structure of small body objects[34, 36, 35, 37, 38, 39, 40], etc. Binary-star systems are important astrophysical objects which are constituted of binary stars revolving around each other. The amount of binary-star systems is huge in the Galaxy. It’s estimated that the amount of binary-star systems is no less than that of single star. The vast majority of stars are located in a binary system. It is estimated that more than 70% of all massive stars may exist in systems with two or three stars.
In active galactic nuclei, due to the rich forms of interactions between SBH, the gaseous accretion disks and the stellar mass compact objects, the binary may be formed and scattered onto the orbits closed to the central SBH to form the systems which we call B-EMRIs. The binary stars orbiting around central SBH constitute a significant typical EMIRs. Thus, it is quite sensible to consider the waveforms produced by motions of binary stars surrounding a SBH in addition to the case of single star. The GW generated by each of the binary star overlaps, so the waveforms generated by this type of EMRIs are distinguishable from that generated by single star. In this paper, we consider the case that the amplitudes of the GW generated by internal motion of binary stars are small compared to that generated by external motion of the mass center since such systems are considerable popular in the Universe and can be detected by the next generation observatory of GWs. It’s expected small perturbations would appear in B-EMRIs when compared to the waveforms of ordinary EMRIs. It’s really worthwhile to build the template waveforms of such systems for future data analysis of space-based detectors for GWs.
@@
Since the traditional method of calculating gravitational waveforms—solving Teukolsky equations—cost a lot of time and computational resources[41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], while realistic data analysis need to handle a large number of waveforms, thus more practical methods are necessary. And it is proved that such methods are feasible. In this paper, we adopt the “numerical kludge” (NK) method which generates waveforms more conveniently and quickly but captures the principle feature of true waveforms[54]. To improve accuracy, aside from the mass quadrupole we take into account the mass octupole and current quadrupole moments of the source to solve the wave equations. It’s justified the NK method does excellently at approximating the true gravitational waveforms so this method has already been used for scoping out data analysis of LISA. For the radiation reaction, we consider the adiabatic evolution of orbit parameters: the timescale of evolution of the orbit parameters is much larger than that of orbit cycles such that the motion of mass point in BH spacetime is nearly geodesic. Further, since the amplitude of GWs generated by internal motions are smaller compared to that generated by external motions of the mass center, to master the principle property of the waveform we concentrate on the radiation reaction due to the GW emissions generated by the motion of mass center. We take the hybrid scheme that combines Tagoshi’s[46], Shibata’s[53] and Ryan’s fluxes[55, 56], whose precision is comparable or higher than Teukolsky-based results for generic circular-inclined orbits [57].
According to equivalence principle only at the origin of free-fall frame (FFF), i.e., the mass center of the binary, the spacetime is flat, if there is departure from the origin the effects of curved spacetime appear. Since for binary system, either star does not locate at the origin of FFF, we should take into account the gravitational effects caused by this reason. Later we’ll study the effects of GEM force on the waveforms of B-EMRIs.
@@ 2 Numerical Kludge Waveforms
The binary are supposed to revolve around each other along elliptic orbit. We introduce the relative coordinates $\vec{\tilde{r}}=\vec{r}_{1}-\vec{r}_{2}$, with $\tilde{r}=\frac{\tilde{p}}{1+\tilde{e}\cos\tilde{\beta}}$, then the coordinates of the binary with masses $m_{1}$ and $m_{2}$ can be expressed with coordinates of mass center and relative coordinates as $\vec{r}_{1}=\vec{r}_{C}+\vec{\tilde{r}}_{1},\vec{r}_{2}=\vec{r}_{C}+\vec{\tilde{r}}_{2}$, where $\vec{\tilde{r}}_{1}=\frac{m_{2}}{m_{1}+m_{2}}\vec{\tilde{r}},\;\vec{\tilde{r}}_{2}=-\frac{m_{1}}{m_{1}+m_{2}}\vec{\tilde{r}}$. We use the coordinates and parameters with tilde to describe the internal motion of binary stars around each other. The spin axis of the central SBH is along $z_{0}$, as displayed in Fig.1, $\theta_{0}$ is the angle between $z_{0}$ and $z$ which is the normal direction of the surface of binary stars’ internal elliptic orbit. $\phi_{0}$ is the angle between $x_{0}$ and the node line which is the intersection line between the surface of binary stars’ internal orbit and the horizonal surface.
Figure 1: A sketch of binary stars orbiting around the central SBH, in which the internal orbit is exaggerated to clearly show the positional relation
@@
The Kerr BH admits two obvious Killing vectors $\partial_{t}$ and $\partial_{\phi}$, which correspond to two conserved charges, energy $E$ and angular momentum $L_{z}$, so for the mass point with mass $\mu$ ($\mu=m_{1}+m_{2}$ in this paper), the action reads
@@
which leads to $p_{r}=dS_{r}/dr$ and $p_{\theta}=dS_{\theta}/d\theta$. Inserting (3) into the Hamilton-Jacobi equation (2) and separating variables we have
@@
Figure 2: The first panel displays the orbit of compact object in EMRIs, the second to fourth panels display the orbits of one of the compact objects in B-EMRIs. The lower four panels display the details of the corresponding orbits in the upper four panels in the x-y plane. For the first panel, the mass of the single compact object is set to be $m=18M_{\odot}$, the initial external orbit parameters are set to be $p(0)=10M,e(0)=0.3,\iota(0)=7\pi/9$. For the second to fourth panels, the masses of the compact objects are set to be $m_{1}=2M_{\odot},m_{2}=3M_{\odot}$; $m_{1}=8M_{\odot},m_{2}=10M_{\odot}$ and $m_{1}=20M_{\odot},m_{2}=30M_{\odot}$ respectively. The initial external orbit parameters are set to be $p(0)=10M,e(0)=0.3$ as well, the internal orbit parameters are set to be $\tilde{p}=0.005M,\tilde{e}=0.2$. The rotation parameter of the SBH is set to be $a=0.9M$.
@@
where $\gamma=E\left[\left(r^{2}+a^{2}\right)^{2}/\Delta-a^{2}\right]-\frac{2MraL_{z}}{\Delta}$. Solving these equations allow us to know the positions of mass center of the binary at any time.
To calculate GWs, we need to calculate the positions of the two compact objects in binary system relative to the mass center. The mass center of the binary system is free-falling around the SBH, thus locally the spacetime around the mass center can be treated as Minkowskian spacetime. Since for binary system the size of their internal orbit is much smaller than that of their external orbit, i.e., $\tilde{p}$ is much smaller than $p$, so the internal motions of binary are dominated by internal Newtonian gravity. For convenience, the two-body problem is reduced to one-body problem, the relative motion is described by the Lagrangian
$L=\frac{1}{2}m_{r}\dot{\vec{\tilde{r}}}^{2}-V^{(i)}(\tilde{r}),$ (13)
where $m_{r}=\frac{m_{1}m_{2}}{m_{1}+m_{2}}$ is the reduced mass, and the potential $V^{(i)}(\tilde{r})=-\frac{m_{1}m_{2}}{\tilde{r}}$ describes the internal gravity.
Solving the aforementioned equations that determine the motions of mass center and the internal motions allow us to give the orbit of the binary. In Fig.2 we show the orbits of one of the compact objects in the binary system. From the lower three detail images in Fig.2 one sees that, due to internal motions the orbit have small fluctuations, and thus the orbits fluctuate more frequently as mass increases. The reason roots in that if keeping other parameters fixed, the cycle of orbit becomes shorter as the mass of gravitational source increases. The fluctuations are expected to give rise to higher-frequency GWs.
@@
Figure 3: $h_{+}$ waveforms of EMRIs and B-EMRIs with different masses. The parameters follow that given in Fig.2, panel by panel.
In the NK prescription, a flat-space trajectory which is “equivalent” to a geodesic in Kerr spacetime is reconstructed by projecting the Boyer-Lindquist coordinates onto a fictitious spherical polar grid. A particle moving along geodesic path in Kerr spacetime can be viewed as the particle moving along curved path in flat-space, so it’s convenient for us to calculate GWs with Cartesian coordinates. In Cartesian coordinates the positions of one of the compact objects is given by
$x_{1}$ $=r\sin\theta\cos\phi-\frac{m_{2}}{m_{1}+m_{2}}\frac{\tilde{p}}{1+\tilde{e}\cos\tilde{\beta}}\left(\sin(\tilde{\beta}+\gamma)\cos\phi_{0}+\cos(\tilde{\beta}+\gamma)\cos\theta_{0}\sin\phi_{0}\right),$ (17)
$y_{1}$ $=r\sin\theta\sin\phi+\frac{m_{2}}{m_{1}+m_{2}}\frac{\tilde{p}}{1+\tilde{e}\cos\tilde{\beta}}\left(-\sin(\tilde{\beta}+\gamma)\sin\phi_{0}+\cos(\tilde{\beta}+\gamma)\cos\theta_{0}\cos\phi_{0}\right),$
$z_{1}$ $=r\cos\theta+\frac{m_{2}}{m_{1}+m_{2}}\frac{\tilde{p}}{1+\tilde{e}\cos\tilde{\beta}}\cos(\tilde{\beta}+\gamma)\sin\theta_{0},$
where $\gamma$ is the angle between the line connecting mass center and one of the compact objects and the line that lies in the plane of binary internal orbit and is perpendicular to the node line. For the position of the other compact object one only has to replace $\frac{m_{2}}{m_{1}+m_{2}}$ with $-\frac{m_{1}}{m_{1}+m_{2}}$ in the above expression.
The expression (15) are valid for a general extended source in flat-space. For a single mass point with mass $m_{1}$ the energy-momentum tensor in flat spacetime is given by
@@
For binary stars we know the energy-momentum tensor is the sum of two such terms with $\mu=m_{1}$ and $m_{2}$ respectively. The momenta (16) now can be simplified to
@@
Substituting (19) into (15) one then obtains the GW generated by a mass point. For binary system the GW is the superposition of that generated by two compact objects.
Figure 4: Comparisons between the $h_{+}$ waveforms of EMRIs and B-EMRIs with distinct semi-latus rectum $\tilde{p}$. The binary masses are fixed to be $m_{1}=8M_{\odot}$ and $m_{2}=10M_{\odot}$, and mass of central SBH is $10^{6}M_{\odot}$. The external orbit parameters are set to be $p(0)=10M,e(0)=0.3,\iota(0)=7\pi/9,a=0.9M$, the internal orbit parameters of the binary stars are set to be $\tilde{e}=0.2,\theta_{0}=\frac{\pi}{3}$ and $\tilde{p}=0.003M,0.004M,0.005M,0.006M,0.007M$ respectively.
@@
where the angles $\{\Theta,\Phi\}$ denote the observation point’s latitude and azimuth respectively. After TT projection the waveform is given by
@@
Figure 5: Comparisons of $h_{+}$ waveforms of EMRIs and B-EMRIs by adjusting $\theta_{0}$. The external orbit parameters are set to be $p(0)=10M,e(0)=0.3,\iota(0)=7\pi/9,a=0.9M$, the internal orbit parameters are set to be $\tilde{p}=0.005M,\tilde{e}=0.2$, and $\theta_{0}=0,\frac{\pi}{6},\frac{\pi}{3},\frac{\pi}{2},\frac{3\pi}{4},\pi$ respectivly.
@@
where $q=a/M$. The angle $\iota$ shows the inclination of internal orbit with respect to the external orbit through the Carter constant,
$Q=L_{z}^{2}\tan^{2}\iota.$ (24)
We defer the expressions of the coefficients $g_{1}(e)\sim g_{14}(e)$ to the Appendix. This expression of fluxes are the results of combining Tagoshi’s[46], Shibata’s [53] and Ryan’s fluxes [55, 56]. Numerical analysis indicates that these results ameliorate the unphysical inspiral properties for nearly circular and polar orbits, such as the inspiral moves rapidly away from circularity for the trajectories near zero eccentricity $e=0$, and there is discontinuity between trajectories with $\iota$ slightly less than $90^{\circ}$ and those slightly above $90^{\circ}$, etc., thus the fluxes (23) produce more physical reasonable results than previous ones. From the fluxes (23), we obtain the evolution of eccentricity and semi-latus rectum
@@
we obtain
$\dot{e}=$ $\frac{1-e^{2}}{2p}\left((1+e)\frac{N_{1}(r_{p})}{N(r_{p})}-(1-e)\frac{N_{1}(r_{a})}{N(r_{a})}\right)\dot{E}+\frac{1-e^{2}}{2p}\left((1+e)\frac{N_{2}(r_{p})}{N(r_{p})}-(1-e)\frac{N_{2}(r_{a})}{N(r_{a})}\right)\dot{L}$ (26)
$+\frac{1-e^{2}}{2p}\left((1+e)\frac{N_{3}(r_{p})}{N(r_{p})}-(1-e)\frac{N_{3}(r_{a})}{N(r_{a})}\right)\dot{\iota},$
$\dot{p}=$ $\left(-\frac{(1+e)^{2}}{2}\frac{N_{1}(r_{p})}{N(r_{p})}-\frac{(1-e)^{2}}{2}\frac{N_{1}(r_{a})}{N(r_{a})}\right)\dot{E}+\left(-\frac{(1+e)^{2}}{2}\frac{N_{2}(r_{p})}{N(r_{p})}-\frac{(1-e)^{2}}{2}\frac{N_{2}(r_{a})}{N(r_{a})}\right)\dot{L}$
$+\left(-\frac{(1+e)^{2}}{2}\frac{N_{3}(r_{p})}{N(r_{p})}-\frac{(1-e)^{2}}{2}\frac{N_{3}(r_{a})}{N(r_{a})}\right)\dot{\iota},$
where we have
$N_{1}(r)$ $=Er^{4}+a^{2}Er^{2}-2aM(L_{z}-aE)r,$ (27)
$N_{2}(r)$ $=-L_{z}\sec^{2}\iota\;r^{2}+2M(L_{z}\sec^{2}\iota-aE)r-a^{2}L_{z}\tan^{2}\iota,$
$N_{3}(r)$ $=L_{z}^{2}\tan\iota\sec^{2}\iota(r(2M-r)-a^{2}),$
$N(r)$ $=-2(\mu^{2}-E^{2})r^{3}+3\mu^{2}Mr^{2}-(a^{2}(\mu^{2}-E^{2})+L_{z}^{2}\sec^{2}\iota)r+L_{z}^{2}M\tan^{2}\iota+(L_{z}-aE)^{2}M.$
Using the relation between inclination angle and Carter constant $Q=L_{z}^{2}\tan^{2}\iota$, one has
$\dot{\iota}=\frac{1}{2L_{z}^{2}\tan(\iota)\sec^{2}(\iota)}\dot{Q}-\frac{\sin(\iota)\cos(\iota)}{L_{z}}\dot{L_{z}}.$ (28)
@@
Figure 6: $h_{+}$ waveforms of EMRIs and B-EMRIs with central BH of mass $10^{8}M_{\odot}$. The masses of the binary are $8M_{\odot}$ and $10M_{\odot}$ (left panel), $30M_{\odot}$ and $30M_{\odot}$ (middle panel), and $80M_{\odot}$ and $100M_{\odot}$ (right panel). The external orbit parameters are set to be $p(0)=10M,e(0)=0.3,\iota(0)=7\pi/9$, the internal motion parameters are set to be $\tilde{p}=0.004M$ and $\tilde{e}=0.2$.